OpenStudy (anonymous):

The slope of the tangent to a curve at any point (x, y) on the curve is -x/y. Find the equation of the curve if the point (2, −2) is on the curve.

OpenStudy (anonymous):

not sure how to do this one

ganeshie8 (ganeshie8):

what do you know about the relation between `slope of tangent line` and `first derivative` ?

OpenStudy (anonymous):

well the first derivative represents the slope of the tangent line on any point of the function you derive

ganeshie8 (ganeshie8):

You're given that the `slope of tangent line` equals -x/y, so setup an equation using that info

OpenStudy (anonymous):

i thought that -x/y was the derivative

ganeshie8 (ganeshie8):

Yes \[\dfrac{dy}{dx}=\dfrac{-x}{y}\]

OpenStudy (anonymous):

what do i solve for though ?

ganeshie8 (ganeshie8):

thats the equation ^

ganeshie8 (ganeshie8):

you need to solve the curve \(y\)

OpenStudy (anonymous):

so implicit differentiation ?

ganeshie8 (ganeshie8):

An usual algebraic equation involves isolating a "variable", but a differential equation involves solving for a "curve"

ganeshie8 (ganeshie8):

familiar with variable separation ?

OpenStudy (anonymous):

yep, give me a sec

OpenStudy (anonymous):

okay im getting \[\sqrt{C - x^2}\]

OpenStudy (anonymous):

is that right?

ganeshie8 (ganeshie8):

looks partially correct, show me ur work

OpenStudy (anonymous):